# Quantum-Enhanced Neural Exchange-Correlation Functionals

### v2

@@ IV.6 Resource estimations for quantum hardware implementation
A robust evaluation of a competitive quantum model ideally involves assessing its time-to-solution, generalization capacity, parameter efficiency, and expressive power. In this context, the resources required for evaluating quantum XC models align with those conventionally utilized for DQC-based QNNs. Kyriienko et al. (2021) We focus on the time-to-solution, which is contingent on the total number of measurements (shots) required for an accurate output evaluation of quantum models. To that end, we examine the number of shots needed for the gradients $\nabla_{\theta}E_{\mathrm{XC}}$, $v_{\mathrm{XC}}[n](\mathbf{r})=$ $\delta E_{\mathrm{XC}}[n]/\delta n(\mathbf{r})$ and energy $E_{\mathrm{XC}}$ evaluations during the training phase.
@@
For a single gradient evaluation, the cost operator $\hat{C}=\sum_{i=1}^{N_{\mathrm{q}}}\hat{Z_{i}},$ in the loss function is diagonal by design. It allows for the simultaneous measurement of all its terms (due to mutual commutation), thereby mitigating associated errors on noisy quantum processors. This drastically reduces the number of shots in comparison to common ab intio variational approaches Peruzzo et al. (2014); McArdle et al. (2019) that require the cost operator to be the electronic Hamiltonian of given system, which typically has many non-commuting terms to be measured (i.e., $\mathcal{O}(N^{4})$ terms, where $N$ is the system size). To evaluate $\nabla_{\theta}E_{\mathrm{XC}}$ or $v_{\mathrm{XC}}$ by using the parameter-shift rule Schuld et al. (2019), two expectation values need to be computed at each epoch during the training (see Appendix D). In practical terms, considering that we have $N_{\text{epochs}}$ during training, we require $N_{\text{shots}}$ per expectation value to achieve the desired accuracy $\epsilon_{\text{exp}}$. Consequently, the total number of shots during training performed for $\nabla_{\theta}E_{\mathrm{XC}}$, taking into account $N_{\text{param}}$ training parameters, amounts to $2N_{\text{shots}}N_{\text{param}}N_{\text{epochs}}$. Similarly for $v_{\mathrm{XC}}$, the gradient $\delta E_{\mathrm{XC}}[n]/\delta n(\mathbf{r})$ is evaluated at $N_{\mathrm{grid}}$ points, resulting in $2N_{\text{shots}}N_{\text{grid}}N_{\text{epochs}}$ shots.
In addition, $E_{\mathrm{XC}}$ needs to be inferred at every KS iteration. In our experience, the number of KS iterations can be set at $N_{\text{KS}}\in[15-100]$ depending on the size and complexity (strongly correlated regime) of the system. The total cost in terms of measurements is given by $(2N_{\text{param}}+2N_{\text{grid}}+N_{\text{KS}})N_{\text{shots}}N_{\text{epochs}}$. We consider the number of trainable parameters $N_{\text{param}}$ that is significantly larger than $N_{\text{KS}}$ but smaller than the grid size $N_{\text{grid}}$, which scales linearly with the number of atoms $N_{\text{grid}}\sim\mathcal{O}(N_{\text{atoms}})$. Hence, the overall cost in terms of measurements is generally expected to scale as $\mathcal{O}(N_{\text{shots}}N_{\text{grid}}N_{\text{epochs}})\sim\mathcal{O}(N_{\mathrm{atoms}})$. This scaling is valid for the global energy embedding (Eq. (5)), a singular energy evaluation suffices ($E_{\text{XC}}[n]\approx F_{\theta}[n]$).
@@
In a local embedding scenario, the XC energy $E_{\text{XC}}$ is approximated through $N_{\text{grid}}$ model evaluations, where the total error is the weighted sum of individual errors, leading to number of shots much greater than with the global energy embedding, $N_{\text{grid}}N_{\text{KS}}N_{\text{shots}}N_{\text{epochs}}\gg N_{\text{KS}}N_{\text{shots}}N_{\text{epochs}}$. Hence, the number of expectation values is $(2N_{\text{param}}+2N_{\text{grid}}+N_{\text{grid}}N_{\text{KS}})N_{\text{shots}}N_{\text{epochs}}$. In terms of scaling with the system size, consider the model fixed (number of parameters is constant) and $N_{\text{grid}}\sim\mathcal{O}(N_{\text{atoms}})$. Then, the number of measurements scales with the system size as $\mathcal{O}(N_{\text{atoms}})$.
@@
Table 5: Scaling of the number of measurements for the training with the proposed local (Eq. (4)) and global (Eq. (5)) energy embeddings in quantum-enhanced KS-DFT.
@@
$E_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{shots}}N_{\text{grid}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{shots}}N_{\text{epochs}})$
$\nabla_{\theta}E_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{shots}}N_{\text{param}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{shots}}N_{\text{param}}N_{\text{epochs}})$
$v_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{shots}}N_{\text{grid}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{shots}}N_{\text{grid}}N_{\text{epochs}})$
Total $\mathcal{O}(N_{\text{grid}})\sim\mathcal{O}(N_{\text{atoms}})$ $\mathcal{O}(N_{\text{grid}})\sim\mathcal{O}(N_{\text{atoms}})$
Thus, the total cost for training with local or global energy embedding scales linearly with the system size in terms of shots due to the need to evaluate the local XC potential on the grid. The aforementioned costs are summarized in Table 5. On quantum hardware, the overall time required for training quantum models depends on the single-shot rate, highly dependent on the hardware type. This consideration ultimately determines the computational resources needed to attain a practical advantage (e.g., lower inference time) over classical XC models.
@@ V Conclusions and outlook
In this work, we developed various (Q)NN-based architectures and evaluated their capabilities to represent XC functionals within KS-DFT. To that end, we extended existing differentiable KS-DFT codes based on JAX and PySCF frameworks, in both 1D and 3D settings, to accommodate quantum models. Next, we defined four classes of models capable of incorporating an arbitrary number of quantum and/or classical layers. As in other works, Li et al. (2021); Kasim and Vinko (2021) we classified the models by their locality (i.e., the fraction of density points considered as the input). To integrate these architectures into KS-DFT, we introduced a global embedding strategy, well suited for quantum architectures due to its minimal measurement requirements on quantum hardware. In terms of scalability, both local and global approaches are influenced by the system size due to the need to compute the local XC potential. However, a significant advantage of our model is that the global energy embedding does not scale with the grid size. This feature substantially improved the accuracy of our model, yielding the best results in our experiments. Finally, we performed numerical simulations of the entire training process on a classical computer.
The results of our experiments demonstrate the promising capabilities of QNNs in efficiently modelling the XC energy functionals and doing so with potentially fewer training parameters due to greater expressivity compared to classical NNs Abbas et al. (2021), with theoretical arguments outlined in Sec. II.3.3. Only fully global (quantum-enhanced) models, where the whole density is considered, produced the best results. In particular, Global QNN provided the simplest, yet sufficiently expressive architecture to model with chemical precision the small molecular systems considered in this work, both in 1D (H2) and 3D (planar H4), even in the presence of strong correlation. With a mere 144 trainable parameters, Global QNN produced similar results to the state-of-the-art Global KSR model, which requires ten times more parameters. The same modest model demonstrated satisfactory generalization capability on the H2-H2 system (two separate dihydrogen molecules), despite being trained solely on H2 and H4. This offers a promising perspective on the potential of QNNs for generalization within this framework.
@@
Additionally, in our QML approach, a quantum-enhanced neural functional can be initially trained and subsequently fine-tuned with additional data as needed. Once trained, this functional can be directly applied to any system to infer its properties. This stands in contrast to variational quantum eigensolver Peruzzo et al. (2014) and imaginary time McArdle et al. (2019) approaches, which necessitate full optimization for each system. The latter methods evolve the quantum state using the exact electronic Hamiltonian. In contrast, QML methods depend on the quality of the training data to ensure qualitatively correct predictions.
@@
As an outlook, in this study, we provided an initial evaluation of proposed techniques on quantum and quantum-inspired neural architectures and molecules. In future works, it would be necessary to expand our training dataset substantially to understand the practical efficiency of quantum-enhanced models. Efforts to construct such datasets are currently ongoing. M Casares et al. (2024) The capability of our models to generalize to significantly larger systems is a critical aspect to investigate. Specifically, can these models effectively learn from small, strongly correlated systems and subsequently generalize accurately to larger systems? Additionally, the FO approach Chai (2012) employed in this study has been evaluated solely on dimers. It remains to be verified whether a single value of the temperature parameter $\gamma$, as detailed in Appendix A, is adequate for training on larger datasets. Alternatively, parameter-free FO approaches, such as the quadratically convergent method can be employed instead Cances et al. (2003). In this work, we opted for the simplest-to-implement method.

### v3

@@ IV.8 Resource estimations for quantum hardware implementation
A robust evaluation of a competitive quantum model ideally involves assessing its time-to-solution, generalization capacity, parameter efficiency, and expressive power. In this context, the resources required for evaluating quantum XC models align with those conventionally utilized for DQC-based QNNs. kyriienko2021solving We focus on the time-to-solution, which is contingent on the total number of measurements (shots) required for an accurate output evaluation of quantum models. To that end, we examine the number of shots $N^{\text{tot}}_{\text{shots}}$ needed for the gradients $\nabla_{\theta}E_{\mathrm{XC}}$, $v_{\mathrm{XC}}[n](\mathbf{r})=$ $\delta E_{\mathrm{XC}}[n]/\delta n(\mathbf{r})$ and energy $E_{\mathrm{XC}}$ evaluations during the training phase.
Assumptions. To reasonably assess the resource requirements, it is important to note that, to achieve a desired precision $\varepsilon$ on the total energy, the required number of epochs $N_{\text{epochs}}$ and the number of model parameters $N_{\text{param}}$ are likely to depend on the problem size, i.e., the number of atoms $N_{\mathrm{atoms}}$. Larger and more diverse datasets may require a more expressive XC model – thus a greater number of parameters – as well as adjustments to the batch size and, consequently, an increase in the number of epochs. In such situations, with large $N_{\text{param}}$ (i.e., large depth), the barren plateau problem may arise, necessitating the use of barren-plateau-resistant architectures schatzki2022theoretical and appropriate mitigation techniques sannia2024engineered. For the sake of providing an approximate estimate, we assume that $N_{\text{epochs}}$ and $N_{\text{param}}$ do not depend on the problem size. The exact scaling would depend on the chosen model and other hyperparameters of the framework. Since neural XC functionals in our approach can be designed to be independent of the problem size, they do not need to change as $N_{\mathrm{atoms}}$ increases, although they may no longer achieve accuracy $\varepsilon$ without architectural adjustment. Thus, our estimates do not guarantee achieving accuracy $\varepsilon$ but instead provide an approximate minimal shot budget under the aforementioned assumptions.
@@
Energy evaluation. For a single evaluation of the XC functional, the cost operator $\hat{C}=\sum_{i=1}^{N_{\mathrm{q}}}\hat{Z_{i}}$, in the loss function is diagonal by design. It allows for the simultaneous measurement of all its terms (due to mutual commutation), thereby mitigating associated errors on noisy quantum processors. This drastically reduces the number of shots in comparison to common ab intio variational approaches peruzzo2014variational; mcardle2019variational that require the cost operator to be the electronic Hamiltonian of given system, which typically has many non-commuting terms to be measured (i.e., $\mathcal{O}(N^{4})$ terms, where $N$ is the system size). In App. E, we establish that $N_{\mathrm{shots}}=\mathcal{O}(\varepsilon^{-2})$ measurements per expectation value are required to achieve an accuracy $\varepsilon$ on the XC energy $E_{\mathrm{XC}}$ when using total magnetization cost operator.
Gradient evaluation. To evaluate $\nabla_{\theta}E_{\mathrm{XC}}$ or $v_{\mathrm{XC}}$ by using the parameter-shift rule schuld2019evaluating, two expectation values need to be computed at each epoch during the training (see Appendix D). In practical terms, considering that we have $N_{\text{epochs}}$ during training, we require $N_{\text{shots}}$ of order $\mathcal{O}(\varepsilon^{-2})$ per expectation value to achieve the desired accuracy $\varepsilon$. Consequently, the total number of shots during training performed for $\nabla_{\theta}E_{\mathrm{XC}}$, taking into account $N_{\text{param}}$ training parameters, amounts to $2N_{\text{shots}}N_{\text{param}}N_{\text{epochs}}$. Similarly for $v_{\mathrm{XC}}$, the gradient $\delta E_{\mathrm{XC}}[n]/\delta n(\mathbf{r})$ is evaluated at $N_{\mathrm{grid}}$ points at $N_{\text{KS}}$ iterations, resulting in $2N_{\text{shots}}N_{\text{KS}}N_{\text{grid}}N_{\text{epochs}}$ shots.
Global embedding. In addition, $E_{\mathrm{XC}}$ needs to be inferred at every KS iteration. In our experience, the number of KS iterations can be set at $N_{\text{KS}}\in[15-100]$ depending on the size and complexity (strongly correlated regime) of the system. See ref. das2023accelerating, where large-scale KS-DFT calculations required up to 30 KS iterations. Therefore, the total cost in terms of measurements in our framework is given by
$N^{\text{tot,g}}_{\text{shots}}=(2N_{\text{param}}+2N_{\text{grid}}N_{\text{KS}}+N_{\text{KS}})N_{\text{shots}}N_{\text{epochs}}.$
For the scaling approximation, we consider the number of trainable parameters $N_{\text{param}}$ and the number of grid points $N_{\text{grid}}$ to be significantly larger than $N_{\text{KS}}$. In the following, we consider $N_{\text{KS}}$ constant. Hence, the overall cost in terms of measurements is generally expected to scale as
$N^{\text{tot,g}}_{\text{shots}}\sim\mathcal{O}((N_{\text{param}}+N_{\text{grid}})N_{\text{shots}}N_{\text{epochs}}).$
This scaling is valid for the global energy embedding (Eq. (5)), where a singular energy evaluation suffices ($E_{\text{XC}}[n]\approx F_{\theta}[n]$).
@@
Local embedding. In a local embedding scenario, the XC energy $E_{\text{XC}}$ is approximated through $N_{\text{grid}}$ model evaluations, where the total error is the weighted sum of individual errors, leading to number of shots much greater than with the global energy embedding, $N_{\text{grid}}N_{\text{KS}}N_{\text{shots}}N_{\text{epochs}}\gg N_{\text{KS}}N_{\text{shots}}N_{\text{epochs}}$. The total number of shots required is
$N^{\text{tot,l}}_{\text{shots}}=(2N_{\text{param}}+2N_{\text{grid}}N_{\text{KS}}+N_{\text{grid}}N_{\text{KS}})N_{\text{shots}}N_{\text{epochs}}.$
Hence, the overall total measurement scaling for local embedding can be estimated as
$N^{\text{tot,l}}_{\text{shots}}\sim\mathcal{O}((N_{\text{param}}+N_{\text{grid}})N_{\text{shots}}N_{\text{epochs}}).$
@@
Table 9: Scaling of the number of measurements for the training with the proposed local (Eq. (4)) and global (Eq. (5)) energy embeddings in quantum-enhanced KS-DFT. We assume that $N_{\text{KS}}$ remains constant das2023accelerating, and define $N^{\text{grid}}_{\text{param}}=N_{\text{grid}}+N_{\text{param}}$.
@@
$E_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{grid}}N_{\text{shots}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{shots}}N_{\text{epochs}})$
$\nabla_{\theta}E_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{param}}N_{\text{shots}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{param}}N_{\text{shots}}N_{\text{epochs}})$
$v_{\mathrm{XC}}$ $\mathcal{O}(N_{\text{grid}}N_{\text{shots}}N_{\text{epochs}})$ $\mathcal{O}(N_{\text{grid}}N_{\text{shots}}N_{\text{epochs}})$
Total $\mathcal{O}(N^{\text{grid}}_{\text{param}}N_{\text{shots}}N_{\text{epochs}})$ $\mathcal{O}(N^{\text{grid}}_{\text{param}}N_{\text{shots}}N_{\text{epochs}})$
In addition, the grid size $N_{\text{grid}}$ scales linearly with the number of atoms $N_{\text{grid}}\sim\mathcal{O}(N_{\text{atoms}})$ becke1988multicenter. Thus, under our assumptions, the total cost for training with local or global energy embedding scales linearly with the system size in terms of shots due to the need to evaluate the local XC potential on the grid. The aforementioned costs are summarized in Table 9.
Concrete estimations. To estimate the shot budget required for an actual implementation on quantum hardware, we assume the typical values used in this work: $N_{\text{epochs}}=10^{3}$, $N_{\text{KS}}=10$, $N_{\text{grid}}=10^{3}$, and $N_{\text{param}}=10^{2}$ (e.g., Global QNN). For the $N_{\mathrm{shots}}$ shots per energy evaluation, Sec. IV.7 shows, for a QNN-based functional, that deterioration from the noiseless result occurs as soon as $N_{\mathrm{shots}}<10^{4}$, consistent with the derivations in App. E. Based on these estimates, achieving $\varepsilon\leq 1$ mHa (chemical precision) is expected to require at least $N_{\mathrm{shots}}\approx 10^{4}-10^{6}$, since the error scales as $\varepsilon=\mathcal{O}(N_{\text{shots}}^{-1/2})$. With the global embedding, the total number of shots is $N^{\text{tot,g}}_{\text{shots}}.$ Hence, per epoch, we require $\approx 10^{8}-10^{10}$ shots, and for the entire training, $N^{\text{tot,g}}_{\text{shots}}\approx 10^{11}-10^{13}$ shots.
The above assumes sequential measurements; however, only the epochs and KS iterations are inherently sequential. The gradients $\nabla_{\theta}E_{\mathrm{XC}}$ and $v_{\mathrm{XC}}$ can be fully parallelized on quantum hardware (i.e., we set $N_{\text{param}}=1$ and $N_{\text{grid}}=1$), reducing the shot requirement to
$N^{\text{tot,g}}_{\text{shots}}\approx N_{\text{KS}}N_{\text{shots}}N_{\text{epochs}},$
which yields $10^{8}-10^{10}$ shots for the entire training. On quantum hardware, the overall time required for training quantum models depends on the single-shot rate, highly dependent on the hardware type. This consideration ultimately determines the computational resources needed to attain a practical advantage (e.g., lower inference time) over classical XC models. Assuming an optimistic measurement time on a superconducting device of 1 ms per shot (e.g., $\approx 1$ kHz shot rate reported in ref. kim2023evidence) and $10^{8}$ total shots, complete training could, in principle, be done in about one day. Additional measurement schemes, such as neural network estimators torlai2020precise, could be employed to reduce variance, which would otherwise prevent achieving chemical precision.
@@ V Conclusions and outlook
In this work, we developed various (Q)NN-based architectures and evaluated their capabilities to represent XC functionals within KS-DFT. To that end, we extended existing differentiable KS-DFT codes based on JAX and PySCF frameworks, in both 1D and 3D settings, to accommodate quantum models. Next, we defined four classes of models capable of incorporating an arbitrary number of quantum and/or classical layers. As in other works, ksr; kasim2021learning we classified the models by their locality (i.e., the fraction of density points considered as the input). To integrate these architectures into KS-DFT, we introduced a global embedding strategy, well suited for quantum architectures due to its minimal measurement requirements on quantum hardware. In terms of scalability, both local and global approaches are influenced by the system size due to the need to compute the local XC potential. However, a significant advantage of our model is that the global energy embedding does not scale with the grid size. This feature substantially improved the accuracy of our model, yielding the best results in our experiments. Finally, we performed numerical simulations of the entire training process on a classical computer.
The results of our experiments demonstrate the promising capabilities of QNNs in efficiently modelling the XC energy functionals. In our proof-of-concept evaluation, we showed that quantum models can achieve performance comparable to classical models with a similar number of parameters (Sec. IV.6), offering small improvements in generalization. Moreover, the potentially greater expressivity of QNNs compared to classical models abbas2021power (e.g., exponential number of Fourier frequencies schuld2021effect as discussed in Sec. II.3.3) may not be strictly necessary for approximating the exact XC functional. Hence, the metrics typically adopted in analyzing quantum ML models (such as expressivity and entangling capability) may not be directly related to the ability of those models to perform well on a specific classical dataset. This observation motivates further investigations on larger datasets and more complex instances of quantum models, a non-trivial task given the inherent challenges of simulating quantum mechanics on classical computers.
Only fully global (quantum-enhanced) models, where the whole density is considered, produced the best results. In particular, Global QNN provided the simplest, yet sufficiently expressive architecture to model with chemical precision the small molecular systems considered in this work, both in 1D (H2) and 3D (planar H4), even in the presence of strong correlation. With a mere 316 trainable parameters, Global QNN produced similar results to the state-of-the-art Global KSR model, which requires ten times more parameters. The same modest model demonstrated satisfactory generalization capability on the H2-H2 system (two separate dihydrogen molecules), despite being trained solely on H2 and H4. This offers a promising perspective on the potential of QNNs for generalization within this framework.
@@
To ensure the stability of our framework in realistic noisy conditions present in current quantum hardware – specifically, the SCF loop, which is the core subroutine of KS-DFT – we developed a theoretical framework in Sec. IV.7 that establishes error bounds and specifies the conditions the model and implementation must satisfy to guarantee well-controlled errors. We verified this theory in a concrete implementation under various noise settings, including typical hardware noise, and recovered the predicted linear scaling. Furthermore, we provided resource estimates indicating the number of measurements required for a hardware implementation, demonstrating the feasibility of our approach when observables are measured in parallel.
@@
Additionally, in our QML approach, a quantum-enhanced neural functional can be initially trained and subsequently fine-tuned with additional data as needed. Once trained, this functional can be directly applied to any system to infer its properties. This stands in contrast to variational quantum eigensolver peruzzo2014variational and imaginary time mcardle2019variational approaches, which necessitate full optimization for each system. The latter methods evolve the quantum state using the exact electronic Hamiltonian. In contrast, QML methods depend on the quality of the training data to ensure qualitatively correct predictions.
@@
As an outlook, in this study, we provided an initial evaluation of proposed techniques on quantum and quantum-inspired neural architectures and molecules. In future works, it would be necessary to expand our training dataset substantially to understand the practical efficiency of quantum-enhanced models. Efforts to construct such datasets are currently ongoing. casares2023graddft The capability of our models to generalize to significantly larger systems is a critical aspect to investigate. Specifically, can these models effectively learn from small, strongly correlated systems and subsequently generalize accurately to larger systems?
Additionally, the FO approach frac-occ employed in this study has been evaluated solely on dimers. It remains to be verified whether a single value of the temperature parameter $\gamma$, as detailed in Appendix A, is adequate for training on larger datasets. Alternatively, parameter-free FO approaches, such as the quadratically convergent method can be employed instead cances2003quadratically. In this work, we opted for the simplest-to-implement method.
An essential direction for future work is to establish rigorous criteria under which QNNs provide genuine advantages over classical models, moving beyond general expressivity arguments. While recent advances in ML-based XC design ksr; kasim2021learning; luise2025accurate have yielded promising results, the quest for a universal functional remains unresolved. To systematically explore this challenge, we have released QEX qex2025pasqal, a flexible framework that supports arbitrary classical-quantum hybrids. Our approach is inherently compatible with barren plateau mitigation techniques (e.g., engineered dissipation sannia2024engineered) and with architectures naturally resilient to trainability issues (e.g., QCNNs cong2019quantum and permutation-equivariant models schatzki2022theoretical). Looking ahead, inspired by quantum-enhanced DFT approaches that leverage QPU-prepared ground states sheridan2024enhancing; koridon2025learning, we propose using QNNs as representation learners that extract task-relevant features from approximate ground-state wavefunctions. Analogous to QCNNs on quantum data cong2019quantum, such models could provide inductive biases tailored to electronic-structure tasks. This perspective positions QNNs not merely as end-to-end regressors for the XC functional, but as modules within a hybrid pipeline for learning more expressive and accurate approximations.
The advent of early fault-tolerant quantum computing (FTQC) architectures is expected to significantly broaden the range of feasible quantum algorithms for chemistry. While fully fledged protocols such as Quantum Phase Estimation (QPE) will ultimately become the method of choice as devices mature, recent studies already show that logical-qubit architectures can outperform variational methods on near-term problems van2024end. Such approaches are fully compatible with our framework and may serve as a natural stepping stone prior to large-scale QPE. In the broader landscape, we anticipate that FTQC-based methods and DFT will coexist as complementary paradigms, with hybrid strategies likely to play a central role in the foreseeable future.